# Adjoint motives of modular forms and the Tamagawa number conjecture

### v1

@@ Abstract.
Let $f$ be a newform of weight $k\geq 2$, level NN with coefficients in a number field $K$, and AA the adjoint motive of the motive $M$ associated to $f$. We carefully discuss the construction of the realisations of $M$ and AA, as well as natural integral structures in these realisations. We then use the method of Taylor and Wiles to verify the $\lambda$-part of the Tamagawa number conjecture of Bloch and Kato for $L(A,0)$ and $L(A,1)$. Here $\lambda$ is any prime of $K$ not dividing $Nk!$, and so that the mod $\lambda$ representation associated to $f$ is absolutely irreducible when restricted to the Galois group over ${\mathbb{Q}}(\sqrt{(-1)^{(\ell-1)/2}\ell})$ where $\lambda\mid\ell$. The method also establishes modularity of all lifts of the mod $\lambda$ representation which are crystalline of Hodge-Tate type (0,k1)(0,k-1).
@@ 0. Introduction
This paper concerns the Tamagawa number conjecture of Bloch and Kato [B-K] for adjoint motives of modular forms of weight $k\geq 2$. The conjecture relates the value at 0 of the associated $L$-function to arithmetic invariants of the motive. We prove that it holds up to powers of certain “bad primes.” The strategy for achieving this is essentially due to Wiles [Wi], as completed with Taylor in [T-W]. The Taylor-Wiles construction yields a formula relating the size of a certain module measuring congruences between modular forms to that of a certain Galois cohomology group. This was carried out in [Wi] and [T-W] in the context of modular forms of weight 22, where it was used to prove results in the direction of the Fontaine-Mazur conjecture [F-M]. While it was no surprise that the method could be generalized to higher weight modular forms and that the resulting formula would be related to the Bloch-Kato conjecture, there remained many technical details to verify in order to accomplish this. In particular, the very formulation of the conjecture relies on a comparison isomorphism between the $\ell$-adic and de Rham realizations of the motive provided by theorems of Faltings [Fa] or Tsuji [Ts], and verification of the conjecture requires the careful application of such a theorem. We also need to generalize results on congruences between modular forms to higher weight, and to compute certain local Tamagawa numbers.
@@ 0.1. Some history
Dirichlet’s class number formula amounts to the conjecture for the Dedekind zeta function for a number field at $s=0$ or 11. The conjecture is also known up to a power of 22 for Dirichlet $L$-functions (including the Riemann zeta function) at any integer ([Ma-W], [B-K], [H-K]). It is known up to an explicit set of bad primes for the $L$-function of a CM elliptic curve at $s=1$ if the order of vanishing is $\leq 1$ ([C-W], [Ru1], [Ko]). There are also partial results for $L$-functions of other modular forms at the central critical value ([G-Z], [Kato3], [Ko-L], [N], [Z]) and for values of certain Hecke $L$-functions ([Hn], [Gu2], [Ki]).
Here we consider the adjoint $L$-function of a modular form of weight $k\geq 2$ at $s=0$ and 11. Special values of the $L$-function associated to the adjoint of a modular form, and more generally, twists of its symmetric square, have been studied by many mathematicians. A method of Rankin relates the values to those of Petersson inner products, and this was used by Ogg [Ogg], Shimura [Sh4], Sturm [St1, St2], Coates and Schmidt [C-S, Schm] to obtain nonvanishing results and rationality results along the lines of Deligne’s conjecture. Hida [Hi1] related the precise value to a number measuring congruences between modular forms. In the case of forms corresponding to (modular) elliptic curves, results relating the value to certain Galois cohomology groups (Selmer groups) were obtained by Coates and Schmidt in the context of Iwasawa theory, and by one of the authors, who in [Fl3] obtained results in the direction of the Bloch-Kato conjecture.
@@ 0.2. The framework
Suppose that $f$ is a newform of weight $k\geq 2$ and level NN. Much of the paper is devoted to the construction of the motivic structure $A_{f}$ for which we prove the conjecture. This construction is not new; it is due for the most part to Eichler, Shimura, Deligne, Jannsen, Scholl and Faltings ([Sh1], [De1], [Ja], [Scho2], [Fa]) We review it however in order to collect the facts we need, provide proofs we could not find in the literature, and set things up in a way suited to the formulation of the Bloch-Kato conjecture.
Let us briefly recall here how the construction works. We start with the modular curve $X_{N}$ parametrizing elliptic curves with level NN structure. Then one takes the Betti, de Rham and $\ell$-adic cohomology of $X_{N}$ with coefficients in a sheaf defined as the (k2)(k-2)-nd symmetric power of the relative cohomology of the universal elliptic curve over $X_{N}$. These come with the additional structure and comparison isomorphisms needed to define a motivic structure $M_{N,k}$, the comparison between $\ell$-adic and de Rham cohomology being provided by a theorem of Faltings [Fa]. The structures $M_{N,k}$ can also be defined as in [Scho2] using Kuga-Sato varieties; this has the advantage of showing they arise from “motives” and provides the option of applying Tsuji’s comparison theorem [Ts]. However the construction using “coefficient sheaves” is better suited to defining and comparing lattices in the realizations which play a key role in the proof.
@@ 0.3. The main theorems
(=Theorem 8.10) Let $f$ be a newform of weight $k\geq 2$ and level NN with coefficients in $K$. If $\lambda$ is not in $T_{f}$, then the $\lambda$-part of the Bloch-Kato conjecture holds for $A_{f}$ and $A_{f}(1)$.
@@
(=Theorem 7.15) Let $f$ be a newform of weight $k\geq 2$ and level NN with coefficients in $K$. Suppose $\Sigma$ is a finite set of rational primes containing those dividing NN. Suppose that $\lambda$ is a prime of $K$ which is not in $T_{f}$ and does not divide any prime in $\Sigma$. Then the ${\mathcal{O}}_{K,\lambda}$-module
@@ 1. Generalities and examples of premotivic structures
Suppose that $F$ is a finite extension of ${\mathbb{Q}}_{p}$ in $\bar{{\mathbb{Q}}}_{p}$. We let $F_{0}$ denote the maximal unramified subextension of $F$ and write $F^{\rm ur}$ for the maximal unramified extension of $F$ in $\bar{{\mathbb{Q}}}_{p}$. Let $\nu$ denote the natural map $G_{F}\to G_{{\mathbb{F}}_{p}}\cong\hat{{\mathbb{Z}}}$ where we view $\hat{{\mathbb{Z}}}$ as being generated by the geometric Frobenius element $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$, which we also view as an automorphism of ${\mathbb{Q}}_{p}^{\rm ur}$. We write $\phi_{p}$ for the arithmetic Frobenius, so $\phi_{p}=\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}^{-1}$. Let $I_{F}$ denote the inertia subgroup of $G_{F}$, $W_{F}$ the Weil subgroup, $\nu^{-1}({\mathbb{Z}})$ and $^{\prime}W_{F}$ the Weil-Deligne group of $F$ (see [De2, §8]). Recall that if $K$ is a field of characteristic zero, then to give a representation of $^{\prime}W_{F}$ with coefficients in $K$ is equivalent to giving a finite-dimensional $K$-vector space $V$ (with the discrete topology), a continuous homomorphism $\rho:W_{F}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{K}V$ and a nilpotent endomorphism NN of $V$ satisfying $\rho(g)N=p^{-\nu(g)}N\rho(g)$. If $F={\mathbb{Q}}_{p}$, we write simply $G_{p}$, $I_{p}$, $W_{p}$ and $^{\prime}W_{p}$ for these groups and identify $I_{p}\subset W_{p}\subset G_{p}$ with their images in $G_{\mathbb{Q}}$.
@@
If $K$ is a number field, then $S_{\bf f}(K)$ denotes the set of finite places of $K$. Suppose that $\lambda\in S_{\bf f}(K)$ divides $\ell\in S_{\bf f}({\mathbb{Q}})$ and $F\subset\bar{{\mathbb{Q}}}_{\ell}$ is a finite extension of $\mathbb{Q}$$_{\ell}$. Let $B_{\mathop{\rm dR}\nolimits}=B_{{\mathop{\rm dR}\nolimits},\ell}$, $B_{\mbox{\scriptsize crys}}=B_{{\mbox{\scriptsize crys}},\ell}$ and $B_{\rm st}=B_{{\rm st},\ell}$ be the rings defined by Fontaine [Fo2, §2],[F-P, I.2.1]. Suppose that $V$ is a finite-dimensional $K_{\lambda}$-representation (i.e., a $\lambda$-adic representation) of $G_{F}$. Then $D_{\mathop{\rm dR}\nolimits}(V)=(B_{\mathop{\rm dR}\nolimits}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is filtered free $F\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank, and $V$ is called de Rham if $\dim_{F}D_{\mathop{\rm dR}\nolimits}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$. Similarly $D_{\mbox{\scriptsize crys}}(V)=(B_{\mbox{\scriptsize crys}}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is a filtered free $F_{0}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank with a $(\phi_{\ell}\otimes 1)$-semilinear endomorphism $\phi$, and $V$ is called crystalline if $\dim_{F_{0}}D_{\mbox{\scriptsize crys}}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$. Similarly $D_{\rm st}(V)=(B_{\rm st}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is a filtered free $F_{0}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank with a $(\phi_{\ell}\otimes 1)$-semilinear endomorphism $\phi$ and an endomorphism NN satisfying $N\phi=\ell\phi N$, and $V$ is called semistable if $\dim_{F_{0}}D_{\rm st}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$. One defines $D_{\mathop{\rm pst}\nolimits}(V)$ as the direct limit of $(B_{\rm st}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F^{\prime}}}$ where $F^{\prime}$ runs over finite extensions of $F$ in $\bar{{\mathbb{Q}}}_{\ell}$; this is a filtered free ${\mathbb{Q}}_{\ell}^{\rm ur}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module and one checks that $(\rho,N)$ with $\rho(g)=g\phi^{\nu(g)}$ makes $D_{\mathop{\rm pst}\nolimits}(V)$ a $K_{\lambda}$-rational representation of $^{\prime}W_{F}$, and $V$ is called potentially semistable if $\dim_{{\mathbb{Q}}_{\ell}^{\rm ur}}D_{\mathop{\rm pst}\nolimits}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$.
@@
  • a decreasing filtration such that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a}A=A$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{b}A=0$ for some $a,b\in{\mathbb{Z}}$, and for each $i\in{\mathbb{Z}}$, $\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}A$ is a direct summand of AA;
@@
It follows from [F-L, 1.8] that ${{\mathcal{M}\mathcal{F}}}({\mathcal{O}}_{F})$ is an abelian category. Let ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ denote the full subcategory of objects AA satisfying $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a}A=A$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+\ell}A=0$ and having no non-trivial quotients $A^{\prime}$ such that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+\ell-1}A^{\prime}=A^{\prime}$, and let ${{\mathcal{M}\mathcal{F}}}^{a}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ denote the full subcategory of ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ consisting of objects of finite length. So ${{\mathcal{M}\mathcal{F}}}^{0}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ is the category denoted $\underline{MF}_{\mathop{\rm tor}\nolimits}^{f,\ell^{\prime}}$ in [F-L], and it follows from [F-L, 6.1] that ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ and ${{\mathcal{M}\mathcal{F}}}^{a}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ are abelian categories, stable under taking subobjects, quotients, direct products and extensions in ${{\mathcal{M}\mathcal{F}}}({\mathcal{O}}_{F})$. We write simply $\mathcal{M}\mathcal{F}$}}, ${{\mathcal{M}\mathcal{F}}}^{a}$, etc. whenever $F={\mathbb{Q}}_{\ell}$.
@@
If AA and $A^{\prime}$ are objects of ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$ such that $A\otimes_{{\mathcal{O}}_{\lambda}}A^{\prime}$ defines an object of ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$, then there is a canonical isomorphism

### v2

@@ Abstract.
Let $f$ be a newform of weight $k\geq 2$, level $N$ with coefficients in a number field $K$, and $A$ the adjoint motive of the motive $M$ associated to $f$. We carefully discuss the construction of the realisations of $M$ and $A$, as well as natural integral structures in these realisations. We then use the method of Taylor and Wiles to verify the $\lambda$-part of the Tamagawa number conjecture of Bloch and Kato for $L(A,0)$ and $L(A,1)$. Here $\lambda$ is any prime of $K$ not dividing $Nk!$, and so that the mod $\lambda$ representation associated to $f$ is absolutely irreducible when restricted to the Galois group over ${\mathbb{Q}}(\sqrt{(-1)^{(\ell-1)/2}\ell})$ where $\lambda\mid\ell$. The method also establishes modularity of all lifts of the mod $\lambda$ representation which are crystalline of Hodge-Tate type (0,k1)(0,k-1).
@@ 0. Introduction
This paper concerns the Tamagawa number conjecture of Bloch and Kato [B-K] for adjoint motives of modular forms of weight $k\geq 2$. The conjecture relates the value at 0 of the associated $L$-function to arithmetic invariants of the motive. We prove that it holds up to powers of certain “bad primes.” The strategy for achieving this is essentially due to Wiles [Wi], as completed with Taylor in [T-W]. The Taylor-Wiles construction yields a formula relating the size of a certain module measuring congruences between modular forms to that of a certain Galois cohomology group. This was carried out in [Wi] and [T-W] in the context of modular forms of weight $2$, where it was used to prove results in the direction of the Fontaine-Mazur conjecture [F-M]. While it was no surprise that the method could be generalized to higher weight modular forms and that the resulting formula would be related to the Bloch-Kato conjecture, there remained many technical details to verify in order to accomplish this. In particular, the very formulation of the conjecture relies on a comparison isomorphism between the $\ell$-adic and de Rham realizations of the motive provided by theorems of Faltings [Fa] or Tsuji [Ts], and verification of the conjecture requires the careful application of such a theorem. We also need to generalize results on congruences between modular forms to higher weight, and to compute certain local Tamagawa numbers.
@@ 0.1. Some history
Dirichlet’s class number formula amounts to the conjecture for the Dedekind zeta function for a number field at $s=0$ or $1$. The conjecture is also known up to a power of $2$ for Dirichlet $L$-functions (including the Riemann zeta function) at any integer ([Ma-W], [B-K], [H-K]). It is known up to an explicit set of bad primes for the $L$-function of a CM elliptic curve at $s=1$ if the order of vanishing is $\leq 1$ ([C-W], [Ru1], [Ko]). There are also partial results for $L$-functions of other modular forms at the central critical value ([G-Z], [Kato3], [Ko-L], [N], [Z]) and for values of certain Hecke $L$-functions ([Hn], [Gu2], [Ki]).
Here we consider the adjoint $L$-function of a modular form of weight $k\geq 2$ at $s=0$ and $1$. Special values of the $L$-function associated to the adjoint of a modular form, and more generally, twists of its symmetric square, have been studied by many mathematicians. A method of Rankin relates the values to those of Petersson inner products, and this was used by Ogg [Ogg], Shimura [Sh4], Sturm [St1, St2], Coates and Schmidt [C-S, Schm] to obtain nonvanishing results and rationality results along the lines of Deligne’s conjecture. Hida [Hi1] related the precise value to a number measuring congruences between modular forms. In the case of forms corresponding to (modular) elliptic curves, results relating the value to certain Galois cohomology groups (Selmer groups) were obtained by Coates and Schmidt in the context of Iwasawa theory, and by one of the authors, who in [Fl3] obtained results in the direction of the Bloch-Kato conjecture.
@@ 0.2. The framework
Suppose that $f$ is a newform of weight $k\geq 2$ and level $N$. Much of the paper is devoted to the construction of the motivic structure $A_{f}$ for which we prove the conjecture. This construction is not new; it is due for the most part to Eichler, Shimura, Deligne, Jannsen, Scholl and Faltings ([Sh1], [De1], [Ja], [Scho2], [Fa]) We review it however in order to collect the facts we need, provide proofs we could not find in the literature, and set things up in a way suited to the formulation of the Bloch-Kato conjecture.
Let us briefly recall here how the construction works. We start with the modular curve $X_{N}$ parametrizing elliptic curves with level $N$ structure. Then one takes the Betti, de Rham and $\ell$-adic cohomology of $X_{N}$ with coefficients in a sheaf defined as the (k2)(k-2)-nd symmetric power of the relative cohomology of the universal elliptic curve over $X_{N}$. These come with the additional structure and comparison isomorphisms needed to define a motivic structure $M_{N,k}$, the comparison between $\ell$-adic and de Rham cohomology being provided by a theorem of Faltings [Fa]. The structures $M_{N,k}$ can also be defined as in [Scho2] using Kuga-Sato varieties; this has the advantage of showing they arise from “motives” and provides the option of applying Tsuji’s comparison theorem [Ts]. However the construction using “coefficient sheaves” is better suited to defining and comparing lattices in the realizations which play a key role in the proof.
@@ 0.3. The main theorems
(=Theorem 8.10) Let $f$ be a newform of weight $k\geq 2$ and level $N$ with coefficients in $K$. If $\lambda$ is not in $T_{f}$, then the $\lambda$-part of the Bloch-Kato conjecture holds for $A_{f}$ and $A_{f}(1)$.
@@
(=Theorem 7.15) Let $f$ be a newform of weight $k\geq 2$ and level $N$ with coefficients in $K$. Suppose $\Sigma$ is a finite set of rational primes containing those dividing $N$. Suppose that $\lambda$ is a prime of $K$ which is not in $T_{f}$ and does not divide any prime in $\Sigma$. Then the ${\mathcal{O}}_{K,\lambda}$-module
@@ 1.1. Galois representations
Suppose that $F$ is a finite extension of ${\mathbb{Q}}_{p}$ in $\bar{{\mathbb{Q}}}_{p}$. We let $F_{0}$ denote the maximal unramified subextension of $F$ and write $F^{\rm ur}$ for the maximal unramified extension of $F$ in $\bar{{\mathbb{Q}}}_{p}$. Let $\nu$ denote the natural map $G_{F}\to G_{{\mathbb{F}}_{p}}\cong\hat{{\mathbb{Z}}}$ where we view $\hat{{\mathbb{Z}}}$ as being generated by the geometric Frobenius element $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$, which we also view as an automorphism of ${\mathbb{Q}}_{p}^{\rm ur}$. We write $\phi_{p}$ for the arithmetic Frobenius, so $\phi_{p}=\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}^{-1}$. Let $I_{F}$ denote the inertia subgroup of $G_{F}$, $W_{F}$ the Weil subgroup, $\nu^{-1}({\mathbb{Z}})$ and $^{\prime}W_{F}$ the Weil-Deligne group of $F$ (see [De2, §8]). Recall that if $K$ is a field of characteristic zero, then to give a representation of $^{\prime}W_{F}$ with coefficients in $K$ is equivalent to giving a finite-dimensional $K$-vector space $V$ (with the discrete topology), a continuous homomorphism $\rho:W_{F}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{K}V$ and a nilpotent endomorphism $N$ of $V$ satisfying $\rho(g)N=p^{-\nu(g)}N\rho(g)$. If $F={\mathbb{Q}}_{p}$, we write simply $G_{p}$, $I_{p}$, $W_{p}$ and $^{\prime}W_{p}$ for these groups and identify $I_{p}\subset W_{p}\subset G_{p}$ with their images in $G_{\mathbb{Q}}$.
@@
If $K$ is a number field, then $S_{\bf f}(K)$ denotes the set of finite places of $K$. Suppose that $\lambda\in S_{\bf f}(K)$ divides $\ell\in S_{\bf f}({\mathbb{Q}})$ and $F\subset\bar{{\mathbb{Q}}}_{\ell}$ is a finite extension of $\mathbb{Q}$$_{\ell}$. Let $B_{\mathop{\rm dR}\nolimits}=B_{{\mathop{\rm dR}\nolimits},\ell}$, $B_{\mbox{\scriptsize crys}}=B_{{\mbox{\scriptsize crys}},\ell}$ and $B_{\rm st}=B_{{\rm st},\ell}$ be the rings defined by Fontaine [Fo2, §2],[F-P, I.2.1]. Suppose that $V$ is a finite-dimensional $K_{\lambda}$-representation (i.e., a $\lambda$-adic representation) of $G_{F}$. Then $D_{\mathop{\rm dR}\nolimits}(V)=(B_{\mathop{\rm dR}\nolimits}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is filtered free $F\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank, and $V$ is called de Rham if $\dim_{F}D_{\mathop{\rm dR}\nolimits}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$. Similarly $D_{\mbox{\scriptsize crys}}(V)=(B_{\mbox{\scriptsize crys}}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is a filtered free $F_{0}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank with a $(\phi_{\ell}\otimes 1)$-semilinear endomorphism $\phi$, and $V$ is called crystalline if $\dim_{F_{0}}D_{\mbox{\scriptsize crys}}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$. Similarly $D_{\rm st}(V)=(B_{\rm st}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is a filtered free $F_{0}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank with a $(\phi_{\ell}\otimes 1)$-semilinear endomorphism $\phi$ and an endomorphism $N$ satisfying $N\phi=\ell\phi N$, and $V$ is called semistable if $\dim_{F_{0}}D_{\rm st}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$. One defines $D_{\mathop{\rm pst}\nolimits}(V)$ as the direct limit of $(B_{\rm st}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F^{\prime}}}$ where $F^{\prime}$ runs over finite extensions of $F$ in $\bar{{\mathbb{Q}}}_{\ell}$; this is a filtered free ${\mathbb{Q}}_{\ell}^{\rm ur}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module and one checks that $(\rho,N)$ with $\rho(g)=g\phi^{\nu(g)}$ makes $D_{\mathop{\rm pst}\nolimits}(V)$ a $K_{\lambda}$-rational representation of $^{\prime}W_{F}$, and $V$ is called potentially semistable if $\dim_{{\mathbb{Q}}_{\ell}^{\rm ur}}D_{\mathop{\rm pst}\nolimits}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$.
@@
  • a decreasing filtration such that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a}A=A$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{b}A=0$ for some $a,b\in{\mathbb{Z}}$, and for each $i\in{\mathbb{Z}}$, $\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}A$ is a direct summand of $A$;
@@
It follows from [F-L, 1.8] that ${{\mathcal{M}\mathcal{F}}}({\mathcal{O}}_{F})$ is an abelian category. Let ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ denote the full subcategory of objects $A$ satisfying $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a}A=A$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+\ell}A=0$ and having no non-trivial quotients $A^{\prime}$ such that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+\ell-1}A^{\prime}=A^{\prime}$, and let ${{\mathcal{M}\mathcal{F}}}^{a}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ denote the full subcategory of ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ consisting of objects of finite length. So ${{\mathcal{M}\mathcal{F}}}^{0}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ is the category denoted $\underline{MF}_{\mathop{\rm tor}\nolimits}^{f,\ell^{\prime}}$ in [F-L], and it follows from [F-L, 6.1] that ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ and ${{\mathcal{M}\mathcal{F}}}^{a}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ are abelian categories, stable under taking subobjects, quotients, direct products and extensions in ${{\mathcal{M}\mathcal{F}}}({\mathcal{O}}_{F})$. We write simply $\mathcal{M}\mathcal{F}$}}, ${{\mathcal{M}\mathcal{F}}}^{a}$, etc. whenever $F={\mathbb{Q}}_{\ell}$.
@@
If $A$ and $A^{\prime}$ are objects of ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$ such that $A\otimes_{{\mathcal{O}}_{\lambda}}A^{\prime}$ defines an object of ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$, then there is a canonical isomorphism